One reaction that produces hydrogen gas can
be represented by the unbalanced chemical
equation
Mg(s) + HCl(aq) −→ MgCl2(aq) + H2(g).
What mass of HCl is consumed by the reaction of 4.45 mol of magnesium?
Answer in units of g.
1
Expert's answer
2019-05-06T07:35:54-0400

Solution: Mg(s) +2HCl(aq) −→ MgCl2(aq) + H2(g)

Amount: 1mol 2mol

Amount: 4.45mol 2×4.45mol

Amount: 24g 2×(1+35.5)g

Mass of HCl consumed =2×4.45mol = 2×4.45×36.5 g=324.85g